The probability of any standard traffic light that you would approach when travelling through town being red is 40%. Given that you hit 5 lights on the way to school, calculate the following:
a. The probability that you hit 0 red lights is
b. The probability that you hit 3 or more red lights is:
c. The expected number of red lights you would hit is:
1
Expert's answer
2020-03-16T13:19:24-0400

There is a binomial distribution with p=0.4 and n=5.


a. "P(X=0)=C^0_n0.4^0(1-0.4)^5=0.0778."


b. "P(X\ge3)=P(X=3)+P(X=4)+P(X=5)="

"C^3_50.4^3(1-0.4)^2+C^4_50.4^4(1-0.4)^1+C^5_50.4^5(1-0.4)^0=0.3174."


c. "E(X)=pn=0.4*5=2."